# Unleashing the Potential of Diffusion Models for Incomplete Data Imputation

### v1

@@ Abstract
This paper introduces DiffPuter, an iterative method for missing data imputation that leverages the Expectation-Maximization (EM) algorithm and Diffusion Models. By treating missing data as hidden variables that can be updated during model training, we frame the missing data imputation task as an EM problem. During the M-step, DiffPuter employs a diffusion model to learn the joint distribution of both the observed and currently estimated missing data. In the E-step, DiffPuter re-estimates the missing data based on the conditional probability given the observed data, utilizing the diffusion model learned in the M-step. Starting with an initial imputation, DiffPuter alternates between the M-step and E-step until convergence. Through this iterative process, DiffPuter progressively refines the complete data distribution, yielding increasingly accurate estimations of the missing data. Our theoretical analysis demonstrates that the unconditional training and conditional sampling processes of the diffusion model align precisely with the objectives of the M-step and E-step, respectively. Empirical evaluations across 10 diverse datasets and comparisons with 16 different imputation methods highlight DiffPuter’s superior performance. Notably, DiffPuter achieves an average improvement of 8.10% in MAE and 5.64% in RMSE compared to the most competitive existing method. The code is available at https://github.com/hengruizhang98/DiffPuter.
@@ 1 Introduction
In the field of data science and machine learning, missing data in tabular datasets is a common issue that can severely impair the performance of predictive models and the reliability of statistical analyses. Missing data can result from various factors, including data entry errors, non-responses in surveys, and system errors during data collection [1, 15, 4]. Properly handling missing data is essential, as improper treatment can lead to biased estimates, reduced statistical power, and invalid conclusions.
@@
A plethora of work proposed over the past decades has propelled the development of missing data imputation research. Early classical methods often relied on partially observed statistical features to impute missing values or were based on shallow machine learning techniques, such as KNN [23], or simple parametric models, such as Bayesian models [17] or Gaussian Mixture Models [5]. These methods offer ample interpretability; however, they are constrained by capacity limitations, making it challenging to achieve satisfying performance. With the advent of deep learning, recent research has primarily focused on predictive [28, 29, 14] or generative deep models [30, 18, 25] for missing data imputation. Predictive models learn to predict the target entries conditioned on other observed entries, guided by masking mechanisms [3] or graph regularization techniques [31, 35]. By contrast, generative methods learn the joint distribution of missing entries and observed entries and try to impute the missing data via conditional sampling [25, 20, 18, 30, 21, 34]. Despite employing state-of-the-art generative models [12, 24, 6], generative-imputation methods still fall short compared to predictive methods. We conjecture that this is due to the following reasons: 1) Incomplete likelihood: generative models need to estimate the joint distribution of missing data and observed data. However, since the missing data itself is unknown, there is an inherent error in the estimated data density. 2) Conditional inference: even if we have a generative model that can generate samples from the complete data distribution faithfully, it is still challenging to perform conditional inference since the learned complete distribution either has very complex parametrization or is learned implicitly, and is hard to perform conditional inference, e.g., conditional sampling.
@@
This paper introduces DiffPuter, a principled generative method for missing data imputation based on the Expectation-Maximization (EM) algorithm [2] and Diffusion Models [27, 10], designed to address the aforementioned issues. The EM algorithm is a well-established method for missing data imputation, capable of addressing the incomplete likelihood issue by iteratively refining the values of the missing data. It operates under the intuitive assumption that knowing the values of the missing variables simplifies the maximum likelihood problem for the imputation task. However, integrating the EM algorithm with deep generative models has been less explored, primarily due to the challenges of performing conditional inference for deep generative models without additional assumptions about distribution parameterization [20, 18, 30]. In this paper, we demonstrate that combining the diffusion model with the EM algorithm creates an effective imputation method for incomplete tabular data. Specifically: 1) In the M-step, DiffPuter employs a diffusion model to learn the joint distribution of the missing and observed data. 2) In the E-step, DiffPuter uses the learned diffusion model to perform flexible and accurate conditional sampling by mixing the forward process for observed entries with the reverse process for missing entries. Theoretically, we show that DiffPuter’s M-step corresponds to the maximum likelihood estimation of the data density, while its E-step represents the Expected A Posteriori (EAP) estimation of the missing values, conditioned on the observed values.
@@
We conducted experiments on 10 benchmark tabular datasets containing both continuous and discrete features under various missing data scenarios. We compared the performance of DiffPuter with 16 competitive imputation methods from different categories. Experimental results demonstrate the superior performance of DiffPuter across all settings and on almost all datasets. In addition, experiments demonstrate that DiffPuter’s iterative training can effectively and gradually reduce the error in density estimation, thereby improving the performance of imputation. Ablation studies also demonstrate DiffPuter’s robustness to different missing mechanisms and missing ratios.
@@ 2 Related Works
Iterative imputation is a widely used approach due to its ability to continuously refine predictions of missing data, resulting in more accurate imputation outcomes. This iterative process is especially crucial for methods requiring an initial estimation of the missing data. The Expectation-Maximization (EM) algorithm [2], a classical method, can be employed for missing data imputation. However, earlier applications often assume simple data distributions, such as mixtures of Gaussians for continuous data or Bernoulli and multinomial densities for discrete data [5]. These assumptions limit the imputation capabilities of these methods due to the restricted density estimation of simple distributions. The integration of EM with deep generative models remains underexplored. A closely related approach is MCFlow [25], which iteratively imputes missing data using normalizing flows [24]. However, MCFlow focuses on recovering missing data through maximum likelihood rather than expectation, and its conditional imputation is achieved through soft regularization instead of precise sampling based on the conditional distribution. Beyond EM, the concept of iterative training is prevalent in state-of-the-art deep learning-based imputation methods. For instance, IGRM [35] constructs a graph from all dataset samples and introduces the concept of friend networks, which are iteratively updated during the imputation learning process. HyperImpute [8] proposes an AutoML imputation method that iteratively refines both model selection and imputed values.
@@
We are not the first to utilize diffusion models for missing data imputation. TabCSDI [34] employs a conditional diffusion model to learn the distribution of masked observed entries conditioned on the unmasked observed entries. MissDiff [21] uses a diffusion model to learn the density of tabular data with missing values by masking the observed entries. Although MissDiff was not originally intended for imputation, it can be easily adapted for this task. Other methods, despite claiming applicability for imputation, are trained on complete data and evaluated on incomplete testing data [32, 9]. These approaches contradict the focus of this study, where the training data itself contains missing values. Additionally, all the aforementioned methods use one-step imputations, which overlook the issue that missing data in the training set can lead to inaccurate data density estimation.
@@ 3.1 Missing Value Imputation for Incomplete Data
We further use $\mathbf{x}^{\rm obs}$ and $\mathbf{x}^{\rm mis}$ to denote the observed data and missing data, respectively (i.e., $\mathbf{x}=(\mathbf{x}^{\rm obs},\mathbf{x}^{\rm mis})$). Note that $\mathbf{x}^{\rm obs}$ is the fixed ground-truth observation, while $\mathbf{x}^{\rm miss}$ is conceptual, unknown, and we we aim to estimate it. The missing value imputation task aims to predict the missing entries $\mathbf{x}^{\rm mis}$ based on the observed entries $\mathbf{x}^{\rm obs}$.
@@
The missing data imputation task can be categorized into two types: in-sample and out-of-sample. In-sample imputation means that the model only has to impute the missing entries in the training set, while out-of-sample imputation requires the model’s capacity to generalize to the unseen data records without fitting its parameters again. Not all imputation methods can generalize to out-of-sample imputation tasks. For example, methods that directly treat the missing values as learnable parameters [19, 33] are hard to apply to unseen records. A desirable imputation method is expected to perform well on both in-sample and out-of-sample imputation tasks, and this paper studies both of the two settings.
@@ 3.2 Formulating Missing Data Imputation with Expectation-Maximization
Treating $\mathbf{x}^{\rm obs}$ as observed variables while $\mathbf{x}^{\rm mis}$ as latent variables, with the estimated density of the complete data distribution being parameterized as $p_{\bm{\theta}}(\mathbf{x})$, we can formulate the missing value imputation problem using the Expectation-Maximization (EM) algorithm. Specifically, when the complete data distribution $p_{\bm{\theta}}(\mathbf{x})$ is available, the optimal estimation of missing values is given by $\mathbf{x}^{{\rm mis}*}={\mathbb{E}}_{\mathbf{x}^{\rm mis}}p(\mathbf{x}^{\rm mis}|\mathbf{x}^{\rm obs},\bm{\theta})$. Conversely, when the missing entries $\mathbf{x}^{\rm mis}$ are known, the density parameters can be optimized via maximum likelihood estimation: $\bm{\theta}^{*}=\mathop{\arg\max}\limits_{\bm{\theta}}p(\mathbf{x}|\bm{\theta})$. Consequently, with the initial estimation of missing values $\mathbf{x}^{\rm mis}$, the model parameters $\bm{\theta}$ and the missing value $\mathbf{x}^{\rm mis}$ can be optimized by iteratively applying M-step and E-step:
@@
Remark 1.
The Expectation step can be equivalently rewritten as $\mathbf{x}^{*}={\mathbb{E}}_{\mathbf{x}}\;p(\mathbf{x}|\mathbf{x}^{\rm obs},\bm{\theta})$, where $\mathbf{x}^{*}=(\mathbf{x}^{{\rm obs}},\mathbf{x}^{{\rm mis}*})$. In the remaining sections, we use $\mathbf{x}^{*}$ to denote the E-step for convenience.
@@ 4 Methodology
In this section, we introduce DiffPuter- Iterative Missing Data Imputation with Diffusion. Based on the Expectation-Maximization (EM) algorithm, DiffPuter updates the density parameter $\bm{\theta}$ and hidden variables $\mathbf{x}^{\rm mis}$ in an iterative manner. Fig. 1 shows the overall architecture and training process of DiffPuter: 1) The M-step fixes the missing entries $\mathbf{x}^{\rm mis}$, then a diffusion model is trained to estimate the density of the complete data distribution $p_{\bm{\theta}}(\mathbf{x})=p_{\bm{\theta}}(\mathbf{x}^{\rm obs},\mathbf{x}^{\rm mis})$; 2) The E-step fixes the model parameters $\bm{\theta}$, then we update the missing entries $\mathbf{x}^{\rm mis}$ via the reverse process of the learned diffusion model $p_{\bm{\theta}}(\mathbf{x})$. The above two steps are executed iteratively until convergence. The following sections introduce the M-step and E-step of DiffPuter, respectively. To avoid confusion, we use $\mathbf{x},\mathbf{x}_{t},\mathbf{x}^{\rm obs}$, etc. to denote samples from real data, $\tilde{\mathbf{x}},\tilde{\mathbf{x}}_{t},\tilde{\mathbf{x}}^{\rm obs}$, etc. to denote samples obtained by the model $\bm{\theta}$, while $\hat{\mathbf{x}},\hat{\mathbf{x}}_{t},\hat{\mathbf{x}}_{\rm obs}$, etc. to denote the specific values of variables.
Figure 1: An overview of the architecture of the proposed DiffPuter. DiffPuter utilizes analog bit encoding to transform discrete variables into continuous ones and use the mean of observed values to initialize the missing entries. The EM algorithm alternates the process of 1) fixing $\mathbf{x}^{\rm mis}$ and estimate diffusion model parameter $\mathbf{\bm{\theta}}$, 2) fixing $\mathbf{\bm{\theta}}$ and estimate $\mathbf{x}^{\rm mis}$, for $K$ iterations. The final imputation result $\mathbf{x}^{*}$ is returned from the E-step of the last iteration.
@@ 4.1 M-step: Density Estimation with Diffusion Models
Given an estimation of complete data $\mathbf{x}=(\mathbf{x}^{\rm obs},\mathbf{x}^{\rm mis})$, M-step aims to learn the density of $\mathbf{x}$, parameterized by model $\bm{\theta}$, i.e., $p_{\bm{\theta}}(\mathbf{x})$. Inspired by the impressive generative modeling capacity of diffusion models [27, 10], DiffPuter learns $p_{\bm{\theta}}(\mathbf{x})$ through a diffusion process, which consists of a forward process that gradually adds Gaussian noises of increasing scales to $\mathbf{x}$, and a reverse process that recovers the clean data from the noisy one:
@@
In the forward process, $\mathbf{x}_{0}=\mathbf{x}=(\mathbf{x}^{\rm obs},\mathbf{x}^{\rm mis})$ is the currently estimated data at time 00, and $\mathbf{x}_{t}$ is the diffused data at time $t$. $\sigma(t)$ is the noise level, i.e., the standard deviation of Gaussian noise, at time $t$. The forward process has defined a series of data distribution $p(\mathbf{x}_{t})=\int_{\mathbf{x}_{0}}p(\mathbf{x}_{t}|\mathbf{x}_{0})p(\mathbf{x}_{0}){\rm d}\mathbf{x}_{0}$, and $p(\mathbf{x}_{0})=p(\mathbf{x})$. Note that when restricting the mean of $\mathbf{x}_{0}$ as 00 and the variance of $\mathbf{x}_{0}$ be small (e.g., via standardization), $p(\mathbf{x}_{t})$ approaches a tractable prior distribution $\pi(\mathbf{x})$ at $t=T$ when $\sigma(T)$ is large enough, meaning $p(\mathbf{x}_{T})\approx\pi(\mathbf{x})$ [26]. In our formulation in Eq. 1, $p(\mathbf{x}_{T})\approx\pi(\mathbf{x})={\mathcal{N}}(\bm{0},\sigma^{2}(T)\mathbf{I})$.
In the reverse process, $\nabla_{\mathbf{x}_{t}}\log p(\mathbf{x}_{t})$ is the gradient of $\mathbf{x}_{t}$’s log-probability w.r.t., to $\mathbf{x}_{t}$, and is also known as the score function. $\bm{\omega}_{t}$ is the standard Wiener process. The model is trained by (conditional) score matching [27], which essentially utilizes a neural network $\bm{\epsilon}_{\bm{\theta}}(\mathbf{x}_{t},t)$ to approximate the conditional score-function $\nabla_{\mathbf{x}_{t}}\log p(\mathbf{x}_{t}|\mathbf{x}_{0})$, which approximates $\nabla_{\mathbf{x}_{t}}\log p(\mathbf{x}_{t})$ in expectation:
@@
Since the score of conditional distribution has analytical solutions, i.e., $\nabla_{\mathbf{x}_{t}}\log p(\mathbf{x}_{t}|\mathbf{x}_{0})=\frac{\nabla_{\mathbf{x}_{t}}p(\mathbf{x}_{t}|\mathbf{x}_{0})}{p(\mathbf{x}_{t}|\mathbf{x}_{0})}=-\frac{\mathbf{x}_{t}-\mathbf{x}_{0}}{\sigma^{2}(t)}=-\frac{\bm{\varepsilon}}{\sigma(t)}$, Eq. 3 can be interpreted as training a neural network $\bm{\epsilon}_{\bm{\theta}}(\mathbf{x}_{t},t)$ to approximate the scaled noise. Therefore, $\bm{\epsilon}_{\bm{\theta}}(\mathbf{x}_{t},t)$ is also known as the denoising function, and in this paper, it is implemented as a five-layer MLP (see Appendix C.4).
@@
(Corollary 1 in [26]) Let $p(\mathbf{x})=p(\mathbf{x}_{0})$ be the data distribution, and $p_{\bm{\theta}}(\mathbf{x})=p_{\bm{\theta}}(\mathbf{x}_{0})$ be the marginal data distribution obtained from the reverse process, then the score-matching loss function in Eq. 3 is an upper bound of the negative-log-likelihood of real data ${\bm{x}}\sim p({\bm{x}})$ over $\bm{\theta}$. Formally,
@@
Remark 3 indicates that the $\bm{\theta}$ optimized by Eq. 3 is the maximum likelihood estimation of data distribution $p(\mathbf{x})$. Consequently, $p_{\bm{\theta}}({\bm{x}})$ approximates accurately $p(\mathbf{x})$ with enough capacity. The algorithmic illustration of DiffPuter’s M-step is summarized in Algorithm 1.

### v2

# DiffPuter: Empowering Diffusion Models for Missing Data Imputation

@@ Abstract
Generative models play an important role in missing data imputation in that they aim to learn the joint distribution of full data. However, applying advanced deep generative models (such as Diffusion models) to missing data imputation is challenging due to 1) the inherent incompleteness of the training data and 2) the difficulty in performing conditional inference from unconditional generative models. To deal with these challenges, this paper introduces DiffPuter, a tailored diffusion model combined with the Expectation-Maximization (EM) algorithm for missing data imputation. DiffPuter iteratively trains a diffusion model to learn the joint distribution of missing and observed data and performs an accurate conditional sampling to update the missing values using a tailored reversed sampling strategy. Our theoretical analysis shows that DiffPuter’s training step corresponds to the maximum likelihood estimation of data density (M-step), and its sampling step represents the Expected A Posteriori estimation of missing values (E-step). Extensive experiments across ten diverse datasets and comparisons with 17 different imputation methods demonstrate DiffPuter’s superior performance. Notably, DiffPuter achieves an average improvement of 6.94% in MAE and 4.78% in RMSE compared to the most competitive existing method.
@@ 1 Introduction
In the field of data science and machine learning, missing data in tabular datasets is a common issue that can severely impair the performance of predictive models and the reliability of statistical analysis. Missing data can result from various factors, including data entry errors, non-responses in surveys, and system errors during data collection (Barnard & Meng, 1999; Lillard et al., 1986; Eckert et al., 2020). Properly handling missing data is essential, as improper treatment can lead to biased estimates, reduced statistical power, and invalid conclusions.
@@
A plethora of work proposed over the past decades has propelled the development of missing data imputation research. Early classical methods often relied on partially observed statistical features to impute missing values or were based on conventional machine learning techniques, such as KNN (Pujianto et al., 2019), or simple parametric models, such as Bayesian models (Ma & Chen, 2018) or Gaussian Mixture Models (García-Laencina et al., 2010). With the advent of deep learning, recent research has primarily focused on predictive (Stekhoven & Bühlmann, 2012; Van Buuren & Karin, 2011; Kyono et al., 2021) or generative deep models (Yoon et al., 2018; Mattei & Frellsen, 2019; Richardson et al., 2020) for missing data imputation. Predictive models learn to predict the target entries conditioned on other observed entries, guided by masking mechanisms (Du et al., 2024) or graph regularization techniques (You et al., 2020; Zhong et al., 2023). By contrast, generative methods learn the joint distribution of missing entries and observed entries and aim to impute the missing data via conditional sampling (Richardson et al., 2020; Nazabal et al., 2020b; Mattei & Frellsen, 2019; Yoon et al., 2018; Ouyang et al., 2023; Zheng & Charoenphakdee, 2022). However, generative imputation methods still fall short compared to predictive methods even when employing state-of-the-art generative models (Kingma & Welling, 2013; Rezende & Mohamed, 2015; Goodfellow et al., 2014; Ho et al., 2020). We think this is primarily due to the incomplete likelihood nature issue in missing data imputation: generative models need to estimate the joint distribution of missing data and observed data. However, since the missing data itself is unknown, there is an inherent error in the estimated data density. The classical Expectation-Maximization (EM) algorithm (Dempster et al., 1977) offers an elegant route to handle this issue, being capable of addressing the incomplete likelihood issue by iteratively refining the values of the missing data.
Integrating EM algorithms with generative models has been extensively studied (Richardson et al., 2020), however, its combination with Diffusion models (Ho et al., 2020), currently the most powerful generative models, is still unexplored. In the M-step, the diffusion models have been shown able to faithfully reconstruct the ground-truth distribution. However, in the E-step, it is usually considered challenging to utilize them to perform condition inference (e.g., predicting missing entries based on observed entries). This is because diffusion models directly model and generate the complete joint distribution of data across all dimensions simultaneously, lacking the flexibility found in VAE and GAN-based approaches (Ma et al., 2020; Peis et al., 2022b; Ma et al., 2018).
@@
This paper introduces DiffPuter, a principled generative method for missing data imputation. DiffPuter explores a path that makes diffusion models compatible with the EM framework, allowing both E-step and M-step to be effective. Specifically: 1) In the M-step, DiffPuter employs a diffusion model to learn the joint distribution of the missing and observed data. With the powerful ability of diffusion models to learn tabular data distributions, in the M-step, DiffPuter learns a parameterized distribution that faithfully recovers the joint distribution of observed and missing entries. 2) In the E-step, DiffPuter uses the learned diffusion model to perform flexible and accurate conditional sampling by mixing the forward process for observed entries with the reverse process for missing entries. Theoretically, we show that DiffPuter’s M-step corresponds to the maximum likelihood estimation of the data density, while its E-step represents the Expected A Posteriori (EAP) estimation of the missing values, conditioned on the observed values.
@@
We conduct experiments [^1] The code is available at https://github.com/hengruizhang98/DiffPuter. on nine benchmark tabular datasets containing both continuous and discrete features under various missing data scenarios. We compare the performance of DiffPuter with 17 competitive imputation methods from different categories. Experimental results demonstrate the superior performance of DiffPuter across all settings and on almost all datasets.
@@ 2 Related Works
Iterative imputation is a widely used approach due to its ability to continuously refine predictions of missing data, resulting in more accurate imputation outcomes. This iterative process is especially crucial for methods requiring an initial estimation of the missing data. The Expectation-Maximization (EM) algorithm (Dempster et al., 1977), a classical method, can be employed for missing data imputation. However, earlier applications often assume simple data distributions, such as mixtures of Gaussians for continuous data or Bernoulli and multinominal densities for discrete data (García-Laencina et al., 2010). These assumptions limit the imputation capabilities of these methods due to the restricted density estimation of simple distributions. The integration of EM with deep generative models remains underexplored. A closely related approach is MCFlow (Richardson et al., 2020), which iteratively imputes missing data using normalizing flows (Rezende & Mohamed, 2015). However, MCFlow focuses on recovering missing data through maximum likelihood rather than expectation, and its conditional imputation is achieved through soft regularization instead of precise sampling based on the conditional distribution. Beyond EM, the concept of iterative training is prevalent in state-of-the-art deep learning-based imputation methods. For instance, IGRM (Zhong et al., 2023) constructs a graph from all dataset samples and introduces the concept of friend networks, which are iteratively updated during the imputation learning process. HyperImpute (Jarrett et al., 2022) proposes an AutoML imputation method that iteratively refines both model selection and imputed values.
@@
We are not the first to utilize diffusion models for missing data imputation. In computer vision, diffusion models have been widely applied in image inpainting, either in the data space (Lugmayr et al., 2022) or the latent space (Corneanu et al., 2024). In the tabular data area, TabCSDI (Zheng & Charoenphakdee, 2022) employs a conditional diffusion model to learn the distribution of masked observed entries conditioned on the unmasked observed entries. MissDiff (Ouyang et al., 2023) uses a diffusion model to learn the density of tabular data with missing values by masking the observed entries. Although MissDiff was not originally intended for imputation, it can be easily adapted for this task. Other methods, despite claiming applicability for imputation, are trained on complete data and evaluated on incomplete testing data (Zhang et al., 2024; Jolicoeur-Martineau et al., 2024). This approach contradicts the focus of this study, where the training data itself contains missing values. Additionally, all the aforementioned methods use one-step imputations, which overlook the issue that missing data in the training set can lead to inaccurate data density estimation. By contrast, the proposed DiffPuter is the first to integrate a diffusion-based generative model into the EM framework. Additionally, we achieved accurate conditional sampling by mixing the forward and reverse processes of diffusion and demonstrated the effectiveness of this approach through theoretical analysis.
@@ 3.1 Missing Value Imputation for Incomplete Data
We further use $\mathbf{x}^{\rm obs}$ and $\mathbf{x}^{\rm mis}$ to denote the observed data and missing data, respectively (i.e., $\mathbf{x}=(\mathbf{x}^{\rm obs},\mathbf{x}^{\rm mis})$). Note that $\mathbf{x}^{\rm obs}$ is the fixed ground-truth observation, while $\mathbf{x}^{\rm mis}$ is conceptual and unknown, and we we aim to estimate it. The missing value imputation task aims to predict the missing entries $\mathbf{x}^{\rm mis}$ based on the observed entries $\mathbf{x}^{\rm obs}$.
@@
The missing data imputation task can be categorized into two types: in-sample and out-of-sample. In-sample imputation means that the model only has to impute the missing entries in the training set, while out-of-sample imputation requires the model’s capacity to generalize to the unseen data records without fitting its parameters again. Not all imputation methods can generalize to out-of-sample imputation tasks. For example, methods that directly treat the missing values as learnable parameters (Muzellec et al., 2020; Zhao et al., 2023) are hard to apply to unseen records. A desirable imputation method is expected to perform well on both in-sample and out-of-sample imputation tasks, and this paper studies both of the two settings.
@@ 3.2 Formulating Missing Data Imputation with Expectation-Maximization
Treating $\mathbf{x}^{\rm obs}$ as observed variables and $\mathbf{x}^{\rm mis}$ as latent variables, with the estimated density of the complete data distribution parameterized as $p_{\bm{\theta}}(\mathbf{x})$, we can formulate the missing value imputation problem using the Expectation-Maximization (EM) algorithm. Specifically, when the complete data distribution $p_{\bm{\theta}}(\mathbf{x})$ is available, the optimal estimation of missing values is given by $\mathbf{x}^{{\rm mis}*}={\mathbb{E}}_{\mathbf{x}^{\rm mis}}p(\mathbf{x}^{\rm mis}|\mathbf{x}^{\rm obs},\bm{\theta})$. Conversely, when the missing entries $\mathbf{x}^{\rm mis}$ are known, the density parameters can be optimized via maximum likelihood estimation: $\bm{\theta}^{*}=\mathop{\arg\max}\limits_{\bm{\theta}}p(\mathbf{x}|\bm{\theta})$. Consequently, with the initial estimation of missing values $\mathbf{x}^{\rm mis}$, the model parameters $\bm{\theta}$ and the missing value $\mathbf{x}^{\rm mis}$ can be optimized by iteratively applying M-step and E-step:
@@ 4 Methodology
In this section, we introduce DiffPuter- Iterative Missing Data Imputation with Diffusion. Based on the Expectation-Maximization (EM) algorithm, DiffPuter updates the density parameter $\bm{\theta}$ and hidden variables $\mathbf{x}^{\rm mis}$ in an iterative manner. Fig. 1 shows the overall architecture and training process of DiffPuter: 1) The M-step fixes the missing entries $\mathbf{x}^{\rm mis}$, then a diffusion model is trained to estimate the density of the complete data distribution $p_{\bm{\theta}}(\mathbf{x})=p_{\bm{\theta}}(\mathbf{x}^{\rm obs},\mathbf{x}^{\rm mis})$; 2) The E-step fixes the model parameters $\bm{\theta}$, then we update the missing entries $\mathbf{x}^{\rm mis}$ via the reverse process of the learned diffusion model $p_{\bm{\theta}}(\mathbf{x})$. The above two steps are executed iteratively until convergence. The following sections introduce the M-step and E-step of DiffPuter, respectively. To avoid confusion, we use $\mathbf{x},\mathbf{x}_{t},\mathbf{x}^{\rm obs}$, etc. to denote samples from real data, $\tilde{\mathbf{x}},\tilde{\mathbf{x}}_{t},\tilde{\mathbf{x}}^{\rm obs}$, etc. to denote samples obtained by the model $\bm{\theta}$, while $\hat{\mathbf{x}},\hat{\mathbf{x}}_{t},\hat{\mathbf{x}}^{\rm obs}$, etc. to denote the specific values of variables.
Figure 1: An overview of the architecture of the proposed DiffPuter. DiffPuter utilizes one-hot encoding to transform discrete variables into continuous ones and use the mean of observed values to initialize the missing entries. The EM algorithm alternates the process of 1) fixing $\mathbf{x}^{\rm mis}$ and estimate diffusion model parameter $\mathbf{\bm{\theta}}$, 2) fixing $\mathbf{\bm{\theta}}$ and estimate $\mathbf{x}^{\rm mis}$, for $K$ iterations. The final imputation result $\mathbf{x}^{*}$ is returned from the E-step of the last iteration.
@@ 4.1 M-step: Density Estimation with Diffusion Models
Given an estimation of complete data $\mathbf{x}=(\mathbf{x}^{\rm obs},\mathbf{x}^{\rm mis})$, M-step aims to learn the density of $\mathbf{x}$, parameterized by model $\bm{\theta}$, i.e., $p_{\bm{\theta}}(\mathbf{x})$. Inspired by the impressive generative modeling capacity of diffusion models (Song et al., 2021b; Karras et al., 2022), DiffPuter learns $p_{\bm{\theta}}(\mathbf{x})$ through a diffusion process, which consists of a forward process that gradually adds Gaussian noises of increasing scales to $\mathbf{x}$, and a reverse process that recovers the clean data from the noisy one:
@@
In the forward process, $\mathbf{x}_{0}=\mathbf{x}=(\mathbf{x}^{\rm obs},\mathbf{x}^{\rm mis})$ is the currently estimated data at time 00, and $\mathbf{x}_{t}$ is the diffused data at time $t$. $\sigma(t)=t$ is the noise level (and $\dot{\sigma}(t)$ is its derivative w.r.t. $t$), i.e., the standard deviation of Gaussian noise, at time $t$. The forward process has defined a series of data distribution $p(\mathbf{x}_{t})=\int_{\mathbf{x}_{0}}p(\mathbf{x}_{t}|\mathbf{x}_{0})p(\mathbf{x}_{0}){\rm d}\mathbf{x}_{0}$, and $p(\mathbf{x}_{0})=p(\mathbf{x})$. Note that when restricting the mean of $\mathbf{x}_{0}$ to 00 and keeping the variance of $\mathbf{x}_{0}$ small (e.g., via standardization), $p(\mathbf{x}_{t})$ approaches a tractable prior distribution $\pi(\mathbf{x})$ at $t=T$ when $\sigma(T)$ is large enough, meaning $p(\mathbf{x}_{T})\approx\pi(\mathbf{x})$ (Song et al., 2021a). In our formulation in Eq. 1, $p(\mathbf{x}_{T})\approx\pi(\mathbf{x})={\mathcal{N}}(\bm{0},\sigma^{2}(T)\mathbf{I})$.
In the reverse process, $\nabla_{\mathbf{x}_{t}}\log p(\mathbf{x}_{t})$ is the gradient of $\mathbf{x}_{t}$’s log-probability w.r.t., to $\mathbf{x}_{t}$, and is also known as the score function. $\bm{\omega}_{t}$ is the standard Wiener process. The model is trained by (conditional) score matching (Song et al., 2021b), which utilizes a neural network $\bm{\epsilon}_{\bm{\theta}}(\mathbf{x}_{t},t)$ (called denoising/score network) to approximate the conditional score-function $\nabla_{\mathbf{x}_{t}}\log p(\mathbf{x}_{t}|\mathbf{x}_{0})$, which in expectation approximates $\nabla_{\mathbf{x}_{t}}\log p(\mathbf{x}_{t})$:
@@
Since the score of conditional distribution has analytical solutions, i.e., $\nabla_{\mathbf{x}_{t}}\log p(\mathbf{x}_{t}|\mathbf{x}_{0})=\frac{\nabla_{\mathbf{x}_{t}}p(\mathbf{x}_{t}|\mathbf{x}_{0})}{p(\mathbf{x}_{t}|\mathbf{x}_{0})}=-\frac{\mathbf{x}_{t}-\mathbf{x}_{0}}{\sigma^{2}(t)}=-\frac{\bm{\varepsilon}}{\sigma(t)}$, Eq. 3 can be interpreted as training a neural network $\bm{\epsilon}_{\bm{\theta}}(\mathbf{x}_{t},t)$ to approximate the scaled noise. Therefore, $\bm{\epsilon}_{\bm{\theta}}(\mathbf{x}_{t},t)$ is also known as the denoising function, and in this paper, it is implemented as a five-layer MLP (see Appendix D.4).
@@
(Corollary 1 in (Song et al., 2021a)) Let $p(\mathbf{x})=p(\mathbf{x}_{0})$ be the data distribution, and $p_{\bm{\theta}}(\mathbf{x})=p_{\bm{\theta}}(\mathbf{x}_{0})$ be the marginal data distribution obtained from the reverse process, then the score-matching loss function in Eq. 3 is an upper bound of the negative-log-likelihood of real data ${\bm{x}}\sim p({\bm{x}})$ over $\bm{\theta}$. Formally,
@@
Remark 2 indicates that the $\bm{\theta}$ optimized by Eq. 3 approximates the maximum likelihood estimation of data distribution $p(\mathbf{x})$. Consequently, $p_{\bm{\theta}}({\bm{x}})$ approximates $p(\mathbf{x})$ accurately with enough capacity. The algorithmic illustration of DiffPuter’s M-step is summarized in Algorithm 1.